OpenStudy (anonymous):

ramona cycles from her house to school at 15mph and returns home at 25mph. if the entire trip takes 32 min, then what is the her avg speed, in miles per hour, for the entire round trip?

OpenStudy (anonymous):

formula for avg. speed??

OpenStudy (anonymous):

not sure but are you referring to d=rt

OpenStudy (anonymous):

It is 5 miles each way: 5 miles / 15 miles per hr = 20 minutes 5 miles / 25 miles per hr = 12 minutes Added, that gives you your 32 minutes. Now, that is 10 miles in 32 minutes = 18.75 mph

OpenStudy (anonymous):

All good now, @lawls ?

OpenStudy (anonymous):

sorry but how'd you know where to start

OpenStudy (anonymous):

(x miles)/(1/4 miles per minute) = 4x = y minutes for the slow leg (x miles)/(5/12 miles per minute) = 12x/5 = 32 - y minutes for the fast leg (32 - y) + y = 32 = 4x + 12x/5 -> x = 5 miles

OpenStudy (anonymous):

The trick is to find the distance.

OpenStudy (anonymous):

okay thank you!

OpenStudy (anonymous):

uw!

OpenStudy (anonymous):

there you go

OpenStudy (anonymous):

uw! Good luck to you in all of your studies and thx for the recognition! @lawls

OpenStudy (anonymous):

no problem and thanks :)